Prove that 2^n>1+ n*√(2^(n-1)) for all n>2 using the inequalities.
1
Expert's answer
2019-03-20T12:05:30-0400

Consider a function and find its derivative:


"f'(n)=(2^{n}-1-n*\sqrt{2^{n-1}})"

Differentiate the sum term by term and factor out constants:


"d(-1)/dn+d(2^n)/dn-d(\sqrt{2^{n-1}}*n)/dn"

The derivative of -1 is zero:


"d(2^n)/dn-d(\sqrt{2^{n-1}}*n)/dn+0"

The derivative of 2^n is 2^nlog(2):


"-(d(\sqrt{2^{n-1}}*n)/dn)+2^nlog(2)"

Use the product rule, "d(u v)\/dn=(v)du\/dn+(u)dv\/dn" where "u=\sqrt{2^{n-1}}" and "v=n"

Simplify the expression:

"2^nlog(2)-n(d(\sqrt{2^{n-1}})/dn)-\sqrt{2^{n-1}}*(d(n)/dn)"

Similarly, we repeat the decomposition and as a result we get:


"-\sqrt{2^{n-1}}+2^n*log(2)-(\sqrt{2^{n-1}}*n*log(2))/2"

Simplify the expression:


"log(2)(2^n-n\sqrt{2^{n-1}})-\sqrt{2^{n-1}}"

As you can see, the derivative is positive when n>2, therefore, the function is increasing at n>2.

Insofar as "f(2)=log(2)(4-\sqrt{2})-\sqrt{2}=0.38" that function positive for all n>2

Thus, for n> 2, the following inequality holds:"f(n)>0, \implies 2^n-1-n\sqrt{2^{n-1}}>0, \implies" "2^n>1+n\sqrt{2^{n-1}}"